. Find the product
3 − i 4
−4 3 + i
! 2 5 − 2i
−5 − 2i 2
!
.
1
Expert's answer
2020-05-28T15:34:53-0400

1) "(3-4i)\times (-43+i)"

"= 3\times-43+3i-4i\times-43-4i^{2}= -125+175i"

Answer : "25(-5+7i)"


2) "(25-2i)\times(-5-22i)"

"= 25 \times -5 + 25 \times -22i - 2i \times -5 + 44i^{2} = -169 -540 i"

Answer :  -169 -540 i